OpenStudy (anonymous):

8(x-2) -5(x+4) = 20+x

OpenStudy (anonymous):

Distribute. 8x-16-5x-20=20+x subtract x 4x-16-20=20 add the 16 and 20 4x=56 divide x=14

OpenStudy (anonymous):

Beat ya to it jazy.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

thanks.. another one coming to ya